Pure carbon reacts with sulfur dioxide gas producing liquid carbon disulfide and carbon monoxide gas. How many liters of carbon monoxide will be produced if 40.0 grams of sulfur dioxide are allowed to react with excess carbon


1
Expert's answer
2021-03-23T04:56:24-0400

The balanced reaction equation:

5C + 2SO2 = CS2 + 4CO

Molecular weight of SO2 is 64 g/mol

The quantity of SO2 is 40 g / (64 g/mol) = 0.625 mol

The quantity of CO is 2 * 0.625 mol = 1.25 mol

The volume of CO is 1.25 mol * (22.4 l/mol) = 28 liters

Answer: 28 litres